OpenStudy (anonymous):

Fractions : 1. Jack spends 1 1/2 as long on his homework than Jill. Last week, Jill spent 8 3/4 hours doing homework. How long did Jack spend doing his homework? 2. And the week before, Jack spent 24 1/2hours, how long did Jill spend doing her homework?

OpenStudy (jack1):

jacks homework time = jills homework time x 1.5 so if jill's homework time = 8 + 3/4 jacks homework time = jills homework time x 1.5 jacks homework time = (8 + 3/4) x 1.5 \[\frac{ 3 }{ 2 }\times (8\frac{ 3 }{ 4 })\]

OpenStudy (missmob):

wouldnt we like convert them into improper first

OpenStudy (anonymous):

When it's x1.5, how would you do that?

OpenStudy (jack1):

1.5 = 1 and 1/2 = 2/2 and 1/2 = 3/2

OpenStudy (anonymous):

What would the answer be?

OpenStudy (jack1):

8 and 3/4 = (8*4)/4 + 3/4 = 32/4 and 3/4 = 35/4

OpenStudy (jack1):

so 3/2 x 35/4 = final answer

OpenStudy (anonymous):

what about part 2?

OpenStudy (jack1):

\[\frac{ 3 }{ 2 }\times \frac{ 35 }{ 4 } = \frac{ 3\times 35 }{ 2 \times 4 } =???\]

OpenStudy (jack1):

part 2: use same equation jacks homework time = jills homework time x 1.5 24.5 = jills homework time x 1.5 24.5 / 1.5 = jills homework time

OpenStudy (anonymous):

so would the first one be, 13 1/8?

OpenStudy (jack1):

yep = 105/8 = 13 and 1/8